OpenStudy (anonymous):

The bank is willing loan you money to purchase a home at 1% interest, compounded monthly. If you would like to buy a $300000 house, hwo much would your payments be each month be if you take out a: 10 year mortgage: 20 year mortgage: 30 year mortgage: Please help!!!

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

formula is rather complicated and long if you want to do this correctly

OpenStudy (anonymous):

derive mortgage formula \begin{align} & 1 \text{st} \text{Month}\text{ }= P(1+R) -W\\&2 \text{nd} \text{Month}\text{ }=((P(1+R) -W)(1+R))-WP(1+R)^2-W(1+R)-W\\&3\text{rd} \text{Month}=\left(\left(P(1+R)^2-W(1+R)-W\right)(1+R)\right)-W\\&P(1+R)^3-W(1+R)^2-W(1+R)-W \end{align}

OpenStudy (anonymous):

\[P(1+R)^N-W\left(-\frac{1-(1+R)^N}{R}\right)=P(1+R)^N+W\frac{\left(1-(1+R)^N\right) }{R}\]

OpenStudy (anonymous):

use bottom formula here with r = .01, y = 10

OpenStudy (anonymous):

\[W\to \frac{P R (1+R)^N}{-1+(1+R)^N}\]

OpenStudy (anonymous):

W= monthly payment P= principal R= interest rate n= number of term

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=300000%28.01%2F12%29%2F%281-%281%2B.01%2F12%29^%28-120%29%29 i get $2826. for the first one

OpenStudy (anonymous):

sorry $2628

OpenStudy (anonymous):

@imran, i don't know that formula. i am wondering if it is the same as the one i wrote

OpenStudy (anonymous):

which formula did you use?

OpenStudy (anonymous):

oops i thought i attached it. bottom one here

OpenStudy (anonymous):

I'm sorry for the late reply but I just wanted to say thank you both so very much for your help with the formulas. I just worked on them and you have given me the correct formula! I used satellite's formula but they both helped me . :)